OpenStudy (anonymous):

The probability that a given urine sample analyzed in a pathology lab contains protein is 0.73 (event A). The probability that the urine sample contains the protein albumin is 0.58 (event B). The probability that a sample contains protein, given that albumin is detected, is 1. Which statement is true? Events A and B are dependent because P(A|B) = P(A). Events A and B are dependent because P(A| B) P(B). Events A and B are dependent because P(A|B) = P(A) + P(B). Events A and B are dependent because P(A| B) P(A). Events A and B are dependent because P(A|B) = P(A).

OpenStudy (amistre64):

dependance has to do with equating a 'given' with a universal probabilty if P(A|B) = P(A|U) then the events are independant

OpenStudy (anonymous):

so its not a or b

OpenStudy (amistre64):

your missing parts of the options so i cant verify that

OpenStudy (amistre64):

The probability that a given urine sample analyzed in a pathology lab contains protein is 0.73 (event A). P(A) = .73 The probability that the urine sample contains the protein albumin is 0.58 (event B). P(B) = .58 The probability that a sample contains protein, given that albumin is detected, is 1. P(AnB) = 1 P(A|B) = P(AnB)/P(B) = 1/.58 does not equal .73 P(A) therefore the events are not independant